Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night report 2019-01-23

Observers: Silvia Piranomonte, Roberta Carini
Support Team: Paolo D'Avanzo, Andrea Melandri
TAT: Maria Teresa Botticella, Enzo Brocato, Lluís Galbany

Daytime Calibrations: Arcs, bias, flat (Gr11,13,16,18,20 - Slit 1.0, 1.5)
Weather:  seeing 1.1

We observed for 3 hours today instead of tomorrow due to a switch time with a DDT program with SOFI (which today had problems with the humidity adapter).
The support astronomer (Emanuela Pompei) asked us to observe today for 3 hours with EFOSC2.
        
UT            OBJECT               SETUP           SLIT          EXP TIME      DIMM     AIRMASS        COMMENTS
02:34       SN2019cj            gr11              1.0            1x2700           1.1''           1.1                  Follow-up
03:36       SN2019cj            gr11              1.0            1x2700           1.5''           1.16                Follow-up
04:40       SN2018jmt          gr13              1.0            1x2700           0.9"          1.20              follow-up   wrong acquisition
04:55       SN2018jmt          gr13              1.0            1x2700           0.9"          1.20              follow-up
05:48       std L745               gr11,gr13     1.0            100,                0.9"          1.10               standard
06:40       SN2019gf            gr13.              1.9            1x1800          1.13"         1.24              follow-up
Comments